The mean cost of domestic airfares in the United States rose to an all-time high of $370 per ticket. Airfares were based on the total ticket value, which consisted of the price charged by the airlines plus any additional taxes and fees. Assume domestic airfares are normally distributed with a standard deviation of $115. Use Table 1 in Appendix B. a. What is the probability that a domestic airfare is $530 or more (to 4 decimals)? 0.0721 b. What is the probability that a domestic airfare is $260 or less (to 4 decimals)? 0.1093 c. What if the probability that a domestic airfare is between $310 and $470 (to 4 decimals)? d. What is the cost for the 3% highest domestic airfares? (rounded to nearest dollar) $ 580 or
Answer #1
  1. What is the probability that a domestic airfare is $530 or more (to 4 decimals)?

µ=$370

α=115

Z =(X- µ)/ α

Z=(530-370)/115 =1.3913

For z=1.3913, probability is= 0.082264 ( from P value table)

So the answer :- The probability that a domestic airfare will cost $530 or more is 0.082264

  1. What is the probability that a domestic airfare is $260 or less (to 4 decimals)?

µ=$370

α=115

Z =(X- µ)/ α

Z=(260-370)/115 =-0.9565

For z=-09565, probability is= 0.16941 ( from P value table)

So the answer :- The probability that a domestic airfare is $260 or less is 0.16941

  1. What if the probability that a domestic airfare is between $310 and $470

For X=310

Z =(X- µ)/ α

Z=(310-370)/115 =-0.5217

For z=-5217, probability is= 0.30094 ( from P value table)

For X=470

Z =(X- µ)/ α

Z=(470-370)/115 =0.8696

For z=0.8696, probability is= 0.19226 ( from P value table)

P(310

Answer= 0.1078